what is the answer to 18y+2=8 as a fraction

Answer by TimothyLamb(4379)

18y + 2 = 8
18y = 8 - 2
18y = 6
y = 6/18
y = 1/3
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---